OpenStudy (anonymous):

in a triangle the second angle is 5 times as large as the first and the third angle in 18 degrees less than 4 times the first angle . find the size of the angles

OpenStudy (mathstudent55):

Let the first angle have measure x. "the second angle is 5 times as large as the first " How would you represent the second angle using x? "the third angle is 18 degrees less than 4 times the first angle" How would you represent the thrid angle in terms of x?

OpenStudy (anonymous):

I did x +5x+4x-18=180 x=19.8 then 5(19.8)=99 how would i do the third

OpenStudy (mathstudent55):

Subtract the sum of the two angle measures you have from 180.

OpenStudy (anonymous):

so is it 19.8-99=180

OpenStudy (anonymous):

can you so me an equationand i work it out

OpenStudy (mathstudent55):

You now know that the first angle is 19.8. The second angle is 99. The sum of all three angles is 180. x + y + z = 180 19.8 + 99 + z = 180 Solve for z, the measure of the third angle.

OpenStudy (anonymous):

61.2?

OpenStudy (mathstudent55):

Correct.